# Limit Theorems For Non-Hermitian Ensembles

### v3

@@ 3.1. Limit theorems of the outer and inner edges of the ring
with $\Gamma(.,.)$ the upper incomplete Gamma function.
A detailed proof of Theorem 7 is provided in Appendix A.2. Numerical results are acknowledged in Figures 4 and 5. The analytical formulation of the probability density function of the scaled minimum modulus $r_{N}$ fits exactly its empirical distribution from the generation of 1010 000000 complex induced Ginibre matrices (left panel) with proportional rectangularity index $L=\alpha N$, $\alpha>0$. As the size of these matrices increases to large numbers (e.g., $N=10e3$), the distribution of $r_{N}$ narrows around the inner radius $r_{in}=1$.
(a) $\alpha=\frac{1}{10}$
(b) $\alpha=\frac{1}{100}(red)\text{, }\alpha=\frac{1}{10}(blue)\text{ and }\alpha=\frac{4}{10}(black)$
Figure 4. (left panel) The Empirical probability distribution (histogram) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=100$, $\alpha=\frac{1}{10}$ and a proportional rectangularity index $L=\alpha N,\alpha>0$. The exact (analytical) probability density function is presented with the solid curve. (right panel) The analytical (exact) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for
$K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=100$. The rectangularity index $L$ is proportional to $N$ such that $L=\alpha N$ with $\alpha>0$. The results are presented for different values of $\alpha=\{\frac{1}{100},\frac{1}{10},\frac{4}{10}\}$. Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
Figure 5. Empirical probability distribution (histogram) and the exact (analytical) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=100$, $\alpha=\frac{1}{10}$ and a proportional rectangularity index $L=\alpha N$. Limiting probability distributions of the scaled minimum modulus $r_{N}$ presented with the red and black curves for large $N=10e4$ and $N=2\times 10e4$, respectively.
Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
The exact formula of the cumulative density function of $r_{N}$ converges towards the asymptotic distribution defined with the location parameter $\mu$ and the scale parameter $\sigma$ of the Gumbel distribution derived from Equation (32), as $N$ goes to infinity (Figure 6).
Figure 6. (left panel) The exact cdf of the scaled minimum modulus $r_{N}$ for $N=10e5$ (blue), $N=10e6$ (blue dotted) and $N=10e7$ (black) and $\alpha=1$. The asymptotic cdf of $r_{N}$ with asymptotic location and scale parameters (red dotted curve). (right panel) The exact formulation of the scaled minimum modulus $r_{N}$ cumulative distribution (black curve) for $N=10e7$ and the asymptotic cumulative distribution (red curve) parametrised with asymptotic location and scale parameters. The parameter $\alpha=1$ for each curve. Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
3.2. Independence of the spectral radius and the minimum of the eigenvalue moduli for the complex induced Ginibre ensemble
@@
Let $\mathcal{B}_{L+N}$ and $\mathcal{B}_{L}$ denote Borel sets in the neighbourhood of $\sqrt{L+N}$ and $\sqrt{L}$, respectively. The events $\{r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}\}$ and $\{r_{min}^{(N)}(G)\in\mathcal{B}_{L}\}$ are independent in the limit as $N$ goes to infinity for fixed rectangularity index $L$.
@@
Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of a $N\times N$ complex induced Ginibre matrix $G$, respectively. The result, presented in Theorem 8, is derived with the formulation of the probability,
@@
where $r$ and $R$ are in $\mathcal{B}_{L}$ and $\mathcal{B}_{L+N}$, respectively. The joint probability density function of the eigenvalues $\left(\lambda_{j}\right)_{j=1}^{N}$ for the complex induced Ginibre ensemble is
$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{N!\pi^{N}}\prod_{k=1}^{N}\frac{1}{\Gamma(k+L)}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)$
Let $r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}$ and $r_{min}^{(N)}(G)\in\mathcal{B}_{L}$ where $\mathcal{B}_{L+N}$ is a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$.
Applying the Andreief’s integration formula, the probability
@@
Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of a $N\times N$ matrices $G$ from the complex induced Ginibre ensemble, respectively. Let $\mathcal{B}_{L+N}$ denote a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$, for proportional rectangularity index $L$. The events $\{r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}\}$ and $\{r_{min}^{(N)}(G)\in\mathcal{B}_{L}\}$ are independent, where $L=\alpha N$ with $\alpha>1$, in the limit as $N$ goes to infinity.
@@
The asymptotics of the regularised lower incomplete Gamma functions is presented in [24]. It is, uniformly for $x\geq 0$, and $a\rightarrow+\infty$ and/or $x\rightarrow+\infty$,
$P(a,x)=\frac{\gamma(a,x)}{\Gamma(a)}=\frac{1}{\Gamma(a)}\int_{0}^{x}t^{a-1}e^{-t}dt=\frac{1}{2}\operatorname{erfc}\left(-\eta(a/2)^{1/2}\right)-\mathcal{R}_{a}(\eta)$
where $\eta=\left[2(\mu-\log(1+\mu))\right]^{1/2}$ with $\mu=\lambda-1$ and $\lambda=\frac{x}{a}$
The remainder $\mathcal{R}_{a}(\eta)$ is represented as
$\mathcal{R}_{a}(\eta)\sim(2\pi a)^{-1/2}e^{-\frac{1}{2}a\eta^{2}}\sum_{k=0}^{+\infty}c_{k}(\eta)^{-k}$
with the coefficients $c_{k}$ defined as in [24] and [25]. These results are (uniformly) valid for $\eta\in\mathbb{R}$.
@@
$\gamma(k+\alpha N,R^{2})=\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)$
@@
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$
$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)}\right]=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}\right)\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$
@@
Furthermore, using the results in [24] and [25] where $\eta=\left[2(\mu-\log(1+\mu))\right]^{1/2}$ with $\mu=\frac{r^{2}}{k+\alpha N}-1$
$\lim_{N\rightarrow+\infty}\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}$ $=\lim_{N\rightarrow+\infty}\left[\frac{1}{2}\operatorname{erfc}\left(-\eta((k+\alpha N)/2)^{1/2}\right)-\mathcal{R}_{t}(\eta)\right]$
$=1-\lim_{N\rightarrow+\infty}\mathcal{R}_{t}(\eta)$
where $t=k+\alpha N$.
The limit of the partial product $\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ is established from the limit of its lower and upper bounds.
Lower bound. The limit
$\lim_{N\rightarrow+\infty}1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}=\lim_{N\rightarrow+\infty}\mathcal{R}_{t}(\eta)$
And for an appropriate choice of $r$, i.e., $r=O\left(\sqrt{\alpha N}\right)$, the parameter $\eta$ is not infinite. This implies that the remainder $\mathcal{R}_{t}(\eta)$ is finite, i.e., $\frac{1}{\mathcal{R}_{t}(\eta)}>C$ where $C$ is a positive constant. This implies that, for $N$ large enough and finite,
$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $>\prod_{k=1}^{N}\left(1+E_{k,N}(R,r)C\right)>\prod_{k=1}^{N}\left(1+\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}(R)C}{N\Gamma(k+\alpha N)^{2}}\right)=1+O\left(\varepsilon_{k,N}(R)\right)$
Upper bound. Using the Stirling formula,
$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $<1+O\left(\frac{\varepsilon_{k,N}(R)e^{-\alpha N}}{\left(N\right)^{\alpha N-\frac{1}{2}-1}}\right)$
@@
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$ $=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}\right]$
$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}$
$=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)=P\left(r_{min}^{(\infty)}(G)\geq r\right)$
@@
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with proportional rectangularity index $L=\alpha N$, $\alpha>0$. The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ are independent random variables, under scaling $\sqrt{\alpha N}$, as $N$ goes to infinity. More precisely, let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$. Setting $\rho=\sqrt{\frac{1+\alpha}{\alpha}}$,
$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\xi_{\alpha}^{(N)}(y)\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\eta_{\alpha}^{(N)}(x)\right]=e^{-e^{y}}e^{-e^{-x}}$
in which $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$ and where
$\xi_{\alpha}^{(N)}(y)=\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }\eta_{\alpha}^{(N)}(x)=\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
The scaled spectral radius $R_{N}$ and the scaled minimum of moduli $r_{N}$ are approximated as
$R_{N}$ $\simeq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }r_{N}\simeq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{Y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
The proof of Theorem 10 is detailed in Appendix A.3.
4. Limiting distribution of the minimum modulus of eigenvalues for matrices from the complex Ginibre and complex induced ensembles. A comparison
The limiting probability distributions of the minimum modulus of eigenvalues for matrices from the complex Ginibre and complex induced Ginibre ensembles are investigated in this section. Derived in Section 2, the survival probability distribution function of the minimum modulus $r_{min}^{(N)}(A)$ of a $N\times N$ complex Ginibre matrix $A$, is used to determine the corresponding analytical probability density function, as $N$ goes infinity. As derived in Section 2,
$p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]$
It represents the exact formulation of the probability density function of the $r_{min}^{(N)}(A)$ as the product of a partial product of the regularised upper incomplete Gamma functions and a partial sum.

### v4

@@ 3.2.1. Limit theorems at the outer and inner edges of the ring
where $\Gamma(.,.)$ is the upper incomplete Gamma function.
Numerical results are acknowledged in Figures 4, 5 and 6. The analytical formulation of the probability density function of the scaled minimum modulus $r_{N}$ fits exactly its empirical distribution from the generation of 1010 000000 complex induced Ginibre matrices (left panel) with proportional rectangularity index $L=\alpha N$, $\alpha>0$. As the size of these matrices increases to large numbers (e.g., $N=10e3$), the distribution of $r_{N}$ narrows around the inner radius $r_{in}=1$.
Figure 4. (left panel) The empirical probability distribution (histogram) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=100$, $\alpha=\frac{1}{10}$ and a proportional rectangularity index $L=\alpha N,\alpha>0$. The exact (analytical) probability density function is presented with the solid curve. (right panel) The analytical (exact) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=100$. The rectangularity index $L$ is proportional to $N$, i.e., $L=\alpha N$ with $\alpha>0$. The results are presented for different values of $\alpha=\{\frac{1}{100},\frac{1}{10},\frac{4}{10}\}$.
Figure 5. Empirical probability distribution (histogram) and the exact (analytical) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=100$, $\alpha=\frac{1}{10}$ and a proportional rectangularity index $L=\alpha N,\forall\alpha>0$. Limiting probability distributions of the scaled minimum modulus $r_{N}$ presented with the red and black curves for large $N=10e4$ and $N=2\times 10e4$, respectively.
The exact formula of the cumulative distribution function of $r_{N}$ converges towards the asymptotic distribution presented in Equation (31), as $N$ goes to infinity (Figure 6).
Figure 6. (left panel) The exact cdf of the scaled minimum modulus $r_{N}$ for $N=10e5$ (blue), $N=10e6$ (blue dotted), $N=10e7$ (black) and $\alpha=1$. The asymptotic cdf of $r_{N}$ (red dotted curve). (right panel) The exact formulation of the scaled minimum modulus $r_{N}$ cumulative distribution function (black curve) for $N=10e7$ and the asymptotic cumulative distribution function (red curve). The parameter $\alpha=1$ for each curve.
3.2.2. Independence of the spectral radius and the minimum modulus for the complex induced Ginibre ensemble
@@
Let $\mathcal{B}_{L+N}$ and $\mathcal{B}_{L}$ denote Borel sets in the neighbourhood of $\sqrt{L+N}$ and $\sqrt{L}$, respectively. Let $r$ and $R$ denote reals in $\mathcal{B}_{L}$ and $\mathcal{B}_{L+N}$, respectively. The events $\{r_{max}^{(N)}(G)\leq R\}$ and $\{r_{min}^{(N)}(G)\geq r\}$ are independent in the limit as $N$ goes to infinity for fixed rectangularity index $L$.
@@
Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of an $N\times N$ complex induced Ginibre matrix $G$, respectively. The result, presented in Theorem 8, is derived with the formulation of the probability,
@@
where $r$ and $R$ are in $\mathcal{B}_{L}$ and $\mathcal{B}_{L+N}$, respectively. The joint probability density function $P_{N}(\lambda_{1},\cdots,\lambda_{N})$ of the eigenvalues $\lambda_{j},j\in\{1,...,N\}$, for the complex induced Ginibre ensemble is the one presented Section 1 with Equation (8).
Let $r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}$ and $r_{min}^{(N)}(G)\in\mathcal{B}_{L}$, where $\mathcal{B}_{L+N}$ is a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$. Applying Andreief’s integration formula, the probability
@@
Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of an $N\times N$ matrix $G$ from the complex induced Ginibre ensemble, respectively. Let $\mathcal{B}_{L+N}$ denote a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$, for proportional rectangularity index $L$. Let $r$ and $R$ denote reals in $\mathcal{B}_{L}$ and $\mathcal{B}_{L+N}$. The events $\{r_{max}^{(N)}(G)\leq R\}$ and $\{r_{min}^{(N)}(G)\geq r\}$ are independent, where $L=\alpha N$ with $\alpha>1$, in the limit as $N$ goes to infinity.
@@
For $R$ in $\mathcal{B}_{L+N}$,
$\gamma(k+\alpha N,R^{2})=\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)$ (33)
@@
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)}\right]$
$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}\right)\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$
@@
The limit of the partial product $\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ is established from the limit of its lower and upper bounds, as $N$ goes to infinity.
Lower bound.
For any fixed $k$ and $N$, $\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}<1$. Let $C$ denote a positive constant such that for any fixed $k$ and $N$, $\frac{1}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}>C$ . This implies that,
$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $>\prod_{k=1}^{N}\left(1+E_{k,N}(R,r)C\right)$
$>\prod_{k=1}^{N}\left(1+\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}(R)C}{N\Gamma(k+\alpha N)^{2}}\right)$
$>\prod_{k=1}^{N}\left(1+\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}(R)C}{N\Gamma(k+\alpha N)\Gamma((1+\alpha)N)}\right)=1+O\left(\frac{\varepsilon_{k,N}(R)}{\Gamma((1+\alpha)N)}\right)$
Upper bound.
$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $<\prod_{k=1}^{N}\left(1+E_{k,N}(R,r)\right)$
$<\prod_{k=1}^{N}\left(1+\frac{\varepsilon_{k,N}(R)}{\Gamma(k+\alpha N)}\right)$
$<\prod_{k=1}^{N}\left(1+\frac{\varepsilon_{k,N}(R)}{(N-1)!}\right)=1+O\left(\frac{\varepsilon_{k,N}(R)}{N^{-1}(N-1)!}\right)$
@@
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$ $=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)$
@@
Theorem 10 (Independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble).
Let $G$ denote an $N\times N$ matrix from the complex induced Ginibre ensemble with proportional rectangularity index $L=\alpha N$, $\alpha=O\left(N^{m}\right),m>1$. The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ are independent random variables, under scaling $\sqrt{\alpha N}$, as $N$ goes to infinity.
More precisely, let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$. Setting $\rho=\sqrt{\frac{1+\alpha}{\alpha}}$,
$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\xi_{\alpha}^{(N)}(y)\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\rho^{2}\alpha N}}+\eta_{\alpha}^{(N)}(x)\right]=e^{-e^{y}}e^{-e^{-x}}$
$\xi_{\alpha}^{(N)}(y)=\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }\eta_{\alpha}^{(N)}(x)=\frac{x}{2\rho\sqrt{2\alpha N\gamma_{\alpha,N}}}$
The scaled spectral radius $R_{N}$ and the scaled minimum of moduli $r_{N}$ are approximated as $N$ goes to infinity, by
$R_{N}$ $\simeq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\rho^{2}\alpha N}}+\frac{X}{2\rho\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }r_{N}\simeq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{Y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
The proof of Theorem 10 is detailed in the Appendix A.3.
3.3. Limiting left and right tail distributions of the minimum modulus for matrices from the complex Ginibre and complex induced ensembles. A comparison.
The limiting left and right tail probability distributions of the minimum modulus for matrices from the complex Ginibre and complex induced Ginibre ensembles are investigated in this section. Derived in Section 3.1, the survival probability distribution function of the minimum modulus $r_{min}^{(N)}(A)$ of an $N\times N$ complex Ginibre matrix $A$, is used to determine the corresponding analytical probability density function.
As derived in Section 3.1,
It represents the exact formulation of the probability density function of the minimum modulus $r_{min}^{(N)}(A)$.
@@
Let $A$ denote an $N\times N$ random matrix from the complex Ginibre ensemble. The limiting left tail probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is the probability density function of the Rayleigh distribution with parameter $\frac{1}{\sqrt{2}}$, as $N$ goes to infinity. More precisely, for $0<r\ll 1$,
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\left(1+O(r^{2})\right)$ (35)
Proof.
The limit of partial product $\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$, as $N$ goes to infinity, has been investigated in Section 3.1 for the complex Ginibre ensemble. For $0<r\ll 1$, it is expressed as
The upper incomplete Gamma function $\Gamma(k,x)$, $k\in\mathbb{N}\setminus\{0\}$, has asymptotic expansion
$\Gamma(k,x)\sim\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$, as $x\longrightarrow 0^{+}$. The limit of the partial sum is derived as follows.
Knowing that, for $0<r\ll 1$,
$\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2(j+1)}\right)$
This implies,
$\lim_{N\rightarrow+\infty}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$ $=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{r^{2j}}{j!+O(r^{2(j+1)})}=2r\left(1+O(r^{2})\right)$
Thus, the limiting left tail probability density function of the minimum modulus for $N\times N$ matrices from the complex Ginibre ensemble is
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)$ $=2re^{-r^{2}}\left(1+O(r^{2})\right)$
This result is in line with the result acknowledged in Theorem 3.
Now, considering any rectangularity index $L\geq 0$, the limiting left tail distribution of the minimum modulus $r_{min}^{(N)}(G)$ for an $N\times N$ random matrix $G$ from the complex induced Ginibre ensemble, as $N$ goes to infinity, is derived as follows.